# Through the Perspective of LiDAR: A Feature-Enriched and Uncertainty-Aware Annotation Pipeline for Terrestrial Point Cloud Segmentation

### v1

@@ 1 Introduction
LiDAR (light detection and ranging) systems are deployed across platforms ranging from satellites to unmanned aerial systems (UAS), mobile units, and terrestrial scanners. Among these, terrestrial laser scanning (TLS) provides exceptionally dense 3D point clouds (mm–cm scale), enabling detailed structural analysis in ecological environments calders2020terrestrial , disney2018weighing . In forest ecosystems, TLS supports the retrieval of tree metrics, biomass, and habitat features liang2016terrestrial , penman2023instructional , and these applications fundamentally rely on accurate semantic segmentation to distinguish ground, stems, branches, foliage, and roots within complex scenes.
While deep learning architectures such as PointNet++ and KPConv have surpassed traditional classifiers (e.g., Random Forests, SVMs) in TLS segmentation qi2017pointnet++ , thomas2019kpconv , their broader utility remains limited by two persistent bottlenecks. First, high-quality annotated datasets are scarce: manual labeling of full-resolution TLS scans is prohibitively labor-intensive, and ecological scenes are especially problematic due to severe occlusion, irregular geometry, and intertwined tree structures kulicki2024artificial , bornand2024completing . Second, the few open-source datasets that do exist are biased toward urban or indoor environments, where objects are simpler and more regularly shaped hackel2017semantic3d , walczak2024inlut3d , dai2017scannet , merizette20253dses . Consequently, current segmentation models often fail to generalize to ecologically complex scenes, slowing the adoption of automated analysis, despite urgent demand in forestry and environmental monitoring.
To overcome these obstacles, we introduce a human-in-the-loop annotation pipeline that reduces labeling burden, while improving segmentation fidelity in challenging forest environments. Our pipeline (i) leverages spherical projection to transform irregular 3D data into structured 2D maps, enabling efficient annotation and feature extraction in a lower-dimensional space; (ii) integrates feature-enriched segmentation to better capture radiometric, geometric, and statistical information; and (iii) incorporates uncertainty analysis to target regions of low model confidence for human correction. Together, these components accelerate annotation, enhance model robustness, and bridge the gap between small, high-effort ecological datasets and scalable, reproducible benchmarks. To this end, we release Mangrove3D—to our knowledge the first TLS benchmark explicitly designed for structurally complex mangrove forests—and demonstrate the pipeline’s cross-domain effectiveness through evaluations on Mangrove3D, ForestSemanticliang2025forestsemantic , mspace_lab_2025_15193973 , and Semantic3Dhackel2017semantic3d .
@@ 1.1.1 Manual Annotation Software and Methodologies
Autodesk ReCapautodesk_recap No Yes No Yes Yes No Engineering, architecture, construction
Trimble RealWorkstrimble_realworks No Yes No Yes Yes No Infrastructure, surveying, civil engineering
Semantic Segmentation Editorsemantic_segmentation_editor Yes Yes Yes Yes Yes No Autonomous driving
LiDAR360semantic_segmentation_editor No Yes Yes Yes Yes Yes Forestry, surveying & mapping, general usage
@@ 1.1.2 Spherical (Equirectangular) Projection for TLS Point Clouds
The concept of spherical projection originates from the equirectangular (plate carrée) projection, which linearly maps longitude and latitude to planar xxyy coordinates. First used in cartography as early as AD 400 snyder1997flattening , this simple yet powerful mapping preserves angular relationships along parallels and meridians. Centuries later, the same principle finds renewed application in LiDAR point‐cloud processing, where unstructured 3D measurements are transformed into structured 2D “range images.” Each point’s azimuth and elevation (or zenith) are linearly mapped to pixel indices, creating a dense angular grid commonly referred to as a spherical projection mao20233d . In essence, this projection provides a way of viewing the world through the perspective of LiDAR—capturing how the sensor perceives its surroundings across its field of view. Scalar attributes such as range, intensity, and return count are then stored as per‐pixel channels, providing a compact 2D representation that preserves the geometric structure of the original 3D scene.
Early works, such as Barnea et al. barnea2008segmentation , BARNEA201333 and Mahmoudabadi et al. MAHMOUDABADI2016135 , leveraged spherical intensity and range maps in combination with classical clustering methods such as mean-shift and graph-based segmentation. These approaches require manual parameter tuning, process each feature channel independently, and are constrained by the limited capacity of traditional machine learning algorithms to integrate heterogeneous features. More recently, DNN-based models such as RangeNet++milioto2019rangenet++ , SalsaNextcortinhal2020salsanext , and SqueezeSegwu2018squeezeseg demonstrated the effectiveness of spherical projection for semantic segmentation, particularly in autonomous driving datasets. However, these approaches were primarily designed for automotive LiDAR, which typically has a narrow vertical field of view and limited feature inputs (often a single channel such as range or intensity), making them strongly dependent on synchronized multi-sensor data.
@@ 1.1.3 Active Learning and Self-Training for Efficient Annotation
Active learning (AL) iteratively queries the annotator for the most informative samples, often those with the highest uncertainty, approximated via Monte-Carlo dropout gal2016dropout or ensembles lakshminarayanan2017simple . While well-studied for image classification, dense segmentation is more difficult, since annotation units (pixel, tile, or superpoint) must balance annotation cost against information gain. Recent studies sener2018active , shao2022active , xu2023hierarchical adapt AL to 2D and 3D domains, but mainly on RGB-rich indoor datasets, and still relies on slow, expertise-intensive 3D point editing.
Self-training is a semi-supervised strategy where a model trained on a seed set reuses its own predictions to expand the labeled pool. Typically, high-confidence predictions are promoted to pseudo-labels and reintroduced for training, sometimes with consistency regularization or noise injection to mitigate overfitting xie2020noisy , zou2020pseudoseg . This approach has shown success in natural images and range-image LiDAR wu2019squeezesegv2 , but faces well-known limitations, e.g., pseudo-label errors easily accumulate, confidence does not always imply correctness, and most pipelines omit human oversight, allowing mistakes to cascade unchecked zou2019confidence , el2025distribution . These challenges are amplified on spherical projection maps of TLS scans, where clutter, occlusion, and class imbalance produce misleading confidence estimates.
We therefore propose an efficient, semi-automatic annotation framework that integrates active learning and self-training within the spherical projection domain to address these challenges. We use an ensemble of UNet++ zhou2018unet++ , DeepLabV3+ chen2018encoder , and Segformer xie2021segformer to quantify pixel-wise epistemic uncertainty, directing annotators to the most informative regions (AL), while high-confidence pixels are automatically promoted to pseudo-labels (self-training). This integrated, uncertainty-aware pipeline reduces manual effort while still ensuring human verification, and— to our knowledge—represents the first scalable framework demonstrated on full-resolution TLS spherical projection maps.
@@ 1.1.4 Feature fusion in semantic segmentation of point clouds
A closely related approach is PointPainting and its derivatives vora2020pointpainting , dong2023pep , which project point clouds onto segmented RGB images, attach semantic scores to points, and feed the enriched clouds into downstream networks. While effective for autonomous driving, these approaches depend on accurately registered multi-sensor data. In contrast, TLS generally operates without co-registered imagery, meaning that spectral cues must be inferred from intrinsic LiDAR signals such as intensity, range, and derived geometric attributes. Our framework is therefore designed to operate solely on LiDAR-derived features, eliminating dependence on external imagery while still capturing complementary structural and radiometric information.
Within LiDAR-only pipelines, feature fusion exploits complementary cues—geometric structure, radiometric response, and positional context—to enhance segmentation. Prior studies have shown that integrating geometric descriptors with intensity and elevation improves class separability chehata2009airborne , rs5083749 , and benchmarks such as Semantic3D confirm gains from combining coordinates, intensity, and shape features hackel2017semantic3d . Accordingly, both projection-based and fully 3D networks (e.g., SqueezeSeg, PointNet++, RandLA-Net) often employ stacked multi-cue inputs qi2017pointnet++ , wu2018squeezeseg , hu2020randla . Yet, despite broad agreement that multi-feature fusion helps, there is still no consensus on an optimal LiDAR-only feature set; redundant or correlated inputs can inflate dimensionality without consistent performance gains weinmann2015semantic .
@@ 2.1 A New TLS Dataset for Mangrove Forest
We use the Canopy Biomass LiDAR (Version 2.0; CBL), a TLS built around a SICK LMS-151 LiDAR unit (SICK AG, Waldkirch, Germany) SICK_LMS151 , which employs a 905 nm wavelength laser and has an effective measurement range of 0.5–50 m at 90% reflectivity. The LiDAR is attached to a rotation stage and the CBL mounted on a modified surface elevation table (SET) arm—a portable leveling device attached to a benchmark pipe, as shown in Fig. 1. For each scan, we invert the CBL from the standard tripod orientation to obtain the maximum potential ground points surrounding the SET, orienting the $90^{\circ}$ uncovered cone above the scanner position, thus enabling $360^{\circ}\times 270^{\circ}$ scans while minimizing occlusion from below. Each CBL scan is acquired at an angular resolution of $0.25^{\circ}$ and completed in 33s.
We collect eight scans at each of the benchmark sites, starting with the first scan position with the SET-arm oriented directly north of the SET and the preceding scans at successive $45^{\circ}$ intervals rotating in a clockwise direction. Scans that suffer from obvious obstruction or operator error are eliminated to ensure data quality.
@@
The dataset comprises 39 TLS scans from seven mangrove plots, totaling 31.3 million points. Individual scans range from 0.51M to 0.94M points (averagely 0.80M per scan), depending on site conditions and preprocessing. Each point is assigned to one of five classes—Ground & Water, Stem, Canopy, Root, and Object—with a sixth label, Void, marking empty pixels in the spherical projection image cube. For segmentation tasks, 30 scans from plots #1–5 are used for training and validation, while 9 scans from plots #6–7 are held out as a fixed test set. In our main experiment, we adopt a train/validation/test split of 27/3/9, yielding a validation-to-train ratio of approximately 0.1. The class distributions across these splits are shown in Fig. 2. A brief annotation guideline is provided in the A.
The Mangrove3D dataset is, to our knowledge, the first TLS benchmark for semantic segmentation in mangrove forests, providing centimeter-scale 3D geometry of tidally influenced ecosystems with dense root and canopy structures. The resulting significant occlusion poses a challenging benchmark for semantic segmentation, a key prerequisite for downstream applications such as biomass estimation and blue-carbon assessments.
@@ 2.2.1 Stage 1: Spherical Projection & Visualization
(a) Point density map
(b) Histogram of the point density
Figure 5: Point‐density validation of the spherical projection map. When the map resolution matches the TLS angular resolution, most pixels contain exactly one point; a value of 0 indicates no return from that angle, 2 indicates a dual‐return measurement, and values above 2 occur rarely, typically due to systematic noise.
@@
Two challenges are scale and density variation. Nearly one million points per scan can lead to excessive memory requirements, and scanner geometry leads to uneven point densities. We partition clouds into azimuth–elevation tiles processed in batches, padded with $k_{\text{b}}$ neighbors to mitigate edge artifacts. For density variation, we use an adaptive neighborhood radius $r_{i}=\text{clamp}(\lambda d_{(k)},r_{\text{min}},r_{\text{max}})$ with $d_{(k)}$ is the $k$-th nearest neighbor distance, $\lambda$ is a scaling factor (e.g., $\lambda=1.5$), and $r_{\text{min}},r_{\text{max}}]=[0.02,0.3]$ m. Together, these strategies enable a scalable, consistent feature computation.
Detailed preprocessing steps, additional feature map examples, correlation analyses, and alternative dimensionality-reduction methods are provided in B (Tables 6, Figs. 21-23).
@@ 2.2.2 Stage 2: Hybrid Annotation with Semi-Supervised and Active Learning
In Stage 2, we adopt a human-in-the-loop strategy that combines self-training and active learning to reduce manual labeling effort, while maintaining high annotation quality (Fig. 6).
A small subset of spherical projection image cubes—one per TLS scan—is manually annotated using Adobe Photoshop (Adobe Inc., 2024) adobe2024 . We use these seed labels for training an ensemble of three diverse 2D semantic segmentation models: UNet++, DeepLabV3+, and Segformer. Each model adopts a unique encoder backbone—i.e., ResNet‑34 he2016deep , EfficientNet‑B3 tan2019efficientnet , and MiT‑B1 yu2023mix , respectively—to promote representational diversity and diminish correlated errors. UNet++ is well-suited for boundary recovery zhou2018unet++ , DeepLabV3+ captures multi-scale context through atrous spatial pyramid pooling chen2018encoder , and SegFormer leverages hierarchical Transformers for long-range dependencies xie2021segformer .
@@
Because the input feature maps extend beyond standard RGB, we re-architect each backbone into a multi-encoder fusion design (Fig. 7(a)). Each 3-channel feature group (e.g., intensity, range, Z-inv) is processed by a dedicated encoder (ResNet-34, EfficientNet-B3, or MiT-B1). These encoders are initialized with ImageNet pre-trained weights, ensuring transfer of learned low-level filters while allowing adaptation to non-RGB feature groups. Deep features are concatenated at the bottleneck and aligned by interpolation before decoding, a strategy shown to outperform naive early fusion rs16203852 , Lei9246289 , ZHANG2020183 . This modular design accommodates any $3\times N$ channel configuration with minimal retraining.
The ensemble combines predictions to generate segmentation masks and epistemic uncertainty maps: high-uncertainty regions are refined manually (active learning), while high-confidence predictions are retained as pseudo-labels (self-training). Fusing logits allows the networks to complement one another dietterich2000ensemble , while deep ensembles provide more reliable uncertainty estimates than a Monte-Carlo dropout approach  lakshminarayanan2017simple , kendall2017uncertainties . The inference–refinement cycle is repeated until all scans are fully annotated. In practice, we display the pseudo-label masks, the uncertainty maps, and the input feature maps to guide annotators. This joint view accelerates review by steering attention to the genuinely ambiguous regions, while leaving well-segmented areas largely untouched. Technical details of the Dice–CrossEntropy joint loss milletari2016v , sudre2017generalised , mao2023cross and the ensemble-based epistemic uncertainty formulation are provided in  C (Eqs. 3-9).

### v3

@@ 1 Introduction
LiDAR (light detection and ranging) systems are deployed across platforms ranging from satellites to unmanned aerial systems (UAS), mobile units, and terrestrial scanners. Among these, terrestrial laser scanning (TLS) provides exceptionally dense 3D point clouds (mm-cm scale), enabling detailed structural analysis in ecological environments Calders et al. [2020], Disney et al. [2018]. In forest ecosystems, TLS supports the retrieval of tree metrics, biomass, and habitat features Liang et al. [2016], Penman et al. [2023], and these applications fundamentally rely on accurate semantic segmentation to distinguish ground, stems, branches, foliage, and roots within complex scenes.
While deep learning architectures such as PointNet++ and KPConv have surpassed traditional classifiers (e.g., Random Forests, SVMs) in TLS segmentation Qi et al. [2017], Thomas et al. [2019], their broader utility remains limited by two persistent bottlenecks. First, high-quality annotated datasets are scarce: manual labeling of full-resolution TLS scans is prohibitively labor-intensive, and ecological scenes are especially problematic due to severe occlusion, irregular geometry, and intertwined tree structures Kulicki et al. [2024], Bornand et al. [2024]. Second, the few open-source datasets that do exist are biased toward urban or indoor environments, where objects are simpler and more regularly shaped Hackel et al. [2017], Walczak [2024], Dai et al. [2017], Mérizette et al. [2025]. Consequently, current segmentation models often fail to generalize to ecologically complex scenes, slowing the adoption of automated analysis, despite urgent demand in forestry and environmental monitoring.
To overcome these obstacles, we introduce a human-in-the-loop annotation pipeline that reduces labeling burden, while improving segmentation fidelity in challenging forest environments. Our pipeline (i) leverages spherical projection to transform irregular 3D data into structured 2D maps, enabling efficient annotation and feature extraction in a lower-dimensional space; (ii) integrates feature-enriched segmentation to better capture radiometric, geometric, and statistical information; and (iii) incorporates uncertainty analysis to target regions of low model confidence for human correction. Together, these components accelerate annotation, enhance model robustness, and bridge the gap between small, high-effort ecological datasets and scalable, reproducible benchmarks. To this end, we release Mangrove3D—to our knowledge the first TLS benchmark explicitly designed for structurally complex mangrove forests—and demonstrate the pipeline’s cross-domain effectiveness through evaluations on Mangrove3D, ForestSemanticLiang et al. [2025], Lab [2025], and Semantic3DHackel et al. [2017].
@@ 1.1.1 Manual Annotation Software and Methodologies
Autodesk ReCapAutodesk [2025] No Yes No Yes Yes No Engineering, architecture, construction
Trimble RealWorksInc. [2025] No Yes No Yes Yes No Infrastructure, surveying, civil engineering
Semantic Segmentation EditorContributors [2025] Yes Yes Yes Yes Yes No Autonomous driving
LiDAR360Contributors [2025] No Yes Yes Yes Yes Yes Forestry, surveying & mapping, general usage
@@ 1.1.2 Spherical (Equirectangular) Projection for TLS Point Clouds
The concept of spherical projection originates from the equirectangular (plate carrée) projection, which linearly maps longitude and latitude to planar $x$-$y$ coordinates. First used in cartography as early as AD 400 Snyder [1997], this simple yet powerful mapping preserves angular relationships along parallels and meridians. Centuries later, the same principle finds renewed application in LiDAR point‐cloud processing, where unstructured 3D measurements are transformed into structured 2D “range images.” Each point’s azimuth and elevation (or zenith) are linearly mapped to pixel indices, creating a dense angular grid commonly referred to as a spherical projection Mao et al. [2023b]. In essence, this projection provides a way of viewing the world through the perspective of LiDAR—capturing how the sensor perceives its surroundings across its field of view. Scalar attributes such as range, intensity, and return count are then stored as per‐pixel channels, providing a compact 2D representation that preserves the geometric structure of the original 3D scene.
Early works, such as Barnea et al. Barnea and Filin [2008, 2013] and Mahmoudabadi et al. Mahmoudabadi et al. [2016], leveraged spherical intensity and range maps in combination with classical clustering methods such as mean-shift and graph-based segmentation. These approaches require manual parameter tuning, process each feature channel independently, and are constrained by the limited capacity of traditional machine learning algorithms to integrate heterogeneous features. More recently, DNN-based models such as RangeNet++Milioto et al. [2019], SalsaNextCortinhal et al. [2020], and SqueezeSegWu et al. [2018] demonstrated the effectiveness of spherical projection for semantic segmentation, particularly in autonomous driving datasets. However, these approaches were primarily designed for automotive LiDAR, which typically has a narrow vertical field of view and limited feature inputs (often a single channel such as range or intensity), making them strongly dependent on synchronized multi-sensor data.
@@ 1.1.3 Active Learning and Self-Training for Efficient Annotation
Active learning (AL) iteratively queries the annotator for the most informative samples, often those with the highest uncertainty, approximated via Monte-Carlo dropout Gal and Ghahramani [2016] or ensembles Lakshminarayanan et al. [2017]. While well-studied for image classification, dense segmentation is more difficult, since annotation units (pixel, tile, or superpoint) must balance annotation cost against information gain. Recent studies Sener and Savarese [2018], Shao et al. [2022], Xu et al. [2023] adapt AL to 2D and 3D domains, but mainly on RGB-rich indoor datasets, and still relies on slow, expertise-intensive 3D point editing.
Self-training is a semi-supervised strategy where a model trained on a seed set reuses its own predictions to expand the labeled pool. Typically, high-confidence predictions are promoted to pseudo-labels and reintroduced for training, sometimes with consistency regularization or noise injection to mitigate overfitting Xie et al. [2020], Zou et al. [2020]. This approach has shown success in natural images and range-image LiDAR Wu et al. [2019], but faces well-known limitations, e.g., pseudo-label errors easily accumulate, confidence does not always imply correctness, and most pipelines omit human oversight, allowing mistakes to cascade unchecked Zou et al. [2019], El Mendili et al. [2025]. These challenges are amplified on spherical projection maps of TLS scans, where clutter, occlusion, and class imbalance produce misleading confidence estimates.
We therefore propose an efficient, semi-automatic annotation framework that integrates active learning and self-training within the spherical projection domain to address these challenges. We use an ensemble of UNet++ Zhou et al. [2018], DeepLabV3+ Chen et al. [2018], and Segformer Xie et al. [2021] to quantify pixel-wise epistemic uncertainty, directing annotators to the most informative regions (AL), while high-confidence pixels are automatically promoted to pseudo-labels (self-training). This integrated, uncertainty-aware pipeline reduces manual effort while still ensuring human verification, and— to our knowledge—represents the first scalable framework demonstrated on full-resolution TLS spherical projection maps.
@@ 1.1.4 Feature fusion in semantic segmentation of point clouds
A closely related approach is PointPainting and its derivatives Vora et al. [2020], Dong et al. [2023], which project point clouds onto segmented RGB images, attach semantic scores to points, and feed the enriched clouds into downstream networks. While effective for autonomous driving, these approaches depend on accurately registered multi-sensor data. In contrast, TLS generally operates without co-registered imagery, meaning that spectral cues must be inferred from intrinsic LiDAR signals such as intensity, range, and derived geometric attributes. Our framework is therefore designed to operate solely on LiDAR-derived features, eliminating dependence on external imagery while still capturing complementary structural and radiometric information.
Within LiDAR-only pipelines, feature fusion exploits complementary cues—geometric structure, radiometric response, and positional context—to enhance segmentation. Prior studies have shown that integrating geometric descriptors with intensity and elevation improves class separability Chehata et al. [2009], Zhang et al. [2013], and benchmarks such as Semantic3D confirm gains from combining coordinates, intensity, and shape features Hackel et al. [2017]. Accordingly, both projection-based and fully 3D networks (e.g., SqueezeSeg, PointNet++, RandLA-Net) often employ stacked multi-cue inputs Qi et al. [2017], Wu et al. [2018], Hu et al. [2020]. Yet, despite broad agreement that multi-feature fusion helps, there is still no consensus on an optimal LiDAR-only feature set; redundant or correlated inputs can inflate dimensionality without consistent performance gains Weinmann et al. [2015].
@@ 2.1 A New TLS Dataset for Mangrove Forest
Table 2: Local surface roughness statistics of the Mangrove3D dataset. Roughness is computed as $\sqrt{\lambda_{3}}$ from local PCA using a radius-based neighborhood (radius = 0.06 m, capped at 60 neighbors) and summarized using robust statistics. Points with insufficient neighbors were excluded from roughness computation.
Label Region Median (cm) Q25–Q75 (cm) P95 (cm) Excluded (NaN) Total points
1 Ground & Water 1.95 1.38–2.59 3.66 62 12,347,797
2 Stem 8.44 5.21–10.37 11.67 1,570 1,788,806
3 Canopy 9.24 7.13–10.59 12.20 49,534 6,096,210
4 Roots 3.43 2.40–6.94 10.71 18,680 10,405,208
5 Objects 2.19 1.67–2.75 3.48 0 373,016
(a) Point density vs range
(b) Point density along height
Figure 2: Point density characteristics across Mangrove3D scans. (a) Point density as a function of range from the scanner, showing individual scans together with the median and interquartile range across scans. (b) Vertical distribution of point returns per scan shown as violin plots, where violin width indicates the relative frequency of points at each height. Scan IDs marked with * correspond to acquisitions with the scanner in an upright orientation; all other scans were acquired with the scanner positioned upside down.
We use the Canopy Biomass LiDAR (Version 2.0; CBL), a TLS built around a SICK LMS-151 LiDAR unit (SICK AG, Waldkirch, Germany) SICK AG [2017], which employs a 905 nm wavelength laser and has an effective measurement range of 0.5–50 m at 90% reflectivity. The LiDAR is attached to a rotation stage and the CBL mounted on a modified surface elevation table (SET) arm—a portable leveling device attached to a benchmark pipe, as shown in Fig. 1. For most scans, we invert the CBL from the standard tripod orientation to obtain the maximum potential ground points surrounding the SET, orienting the $90^{\circ}$ uncovered cone above the scanner position, thus enabling $360^{\circ}\times 270^{\circ}$ scans while minimizing occlusion from below. Fig. 2 provides a compact validation of point density characteristics across all scans. Density decreases consistently with range across scans, while the vertical distributions confirm a strong near-ground concentration of points, consistent with the root-focused acquisition strategy. To assess geometric noise characteristics, Table 2 summarizes PCA-based local surface roughness statistics of the Mangrove3D dataset across major semantic regions. Lower roughness is observed for ground and root surfaces, while higher variability is present in stem and canopy regions, consistent with reduced point density at larger ranges.
Each CBL scan is acquired at an angular resolution of $0.25^{\circ}$ and completed in 33s. We collect eight scans at each of the benchmark sites, starting with the first scan position with the SET-arm oriented directly north of the SET and the preceding scans at successive $45^{\circ}$ intervals rotating in a clockwise direction. Scans that suffer from obvious obstruction or operator error are eliminated to ensure data quality.
@@
The dataset comprises 39 TLS scans from seven mangrove plots, totaling 31.3 million points. Individual scans range from 0.51M to 0.94M points (averagely 0.80M per scan), depending on site conditions and preprocessing. Each point is assigned to one of five classes—Ground & Water, Stem, Canopy, Root, and Object—with a sixth label, Void, marking empty pixels in the spherical projection image cube. For segmentation tasks, 30 scans from plots #1-5 are used for training and validation, while 9 scans from plots #6-7 are held out as a fixed test set. In our main experiment, we adopt a train/validation/test split of 27/3/9, yielding a validation-to-train ratio of approximately 0.1. The class distributions across these splits are shown in Fig. 3. A brief annotation guideline is provided in the A. The Mangrove3D dataset is, to our knowledge, the first TLS benchmark for semantic segmentation in mangrove forests, providing centimeter-scale 3D geometry of tidally influenced ecosystems with dense root and canopy structures. The resulting significant occlusion poses a challenging benchmark for semantic segmentation, a key prerequisite for downstream applications such as biomass estimation and blue-carbon assessments.
@@ 2.2.1 Stage 1: Spherical Projection & Visualization
(a) Spherical projection point-density map with corresponding percentage distribution for a representative scan (same as in Figure 5.)
(b) Scan-wise point-density distribution across the Mangrove3D dataset (39 scans).
Figure 6: Point-density characteristics of the spherical projection map. When the map resolution matches the TLS angular resolution, 80% - 90% of pixels contain exactly one point. Pixels with a value of 0 indicate no return from the corresponding angular direction, values of 2 represent dual-return measurements (mostly <10%), and values greater than 2 occur rarely, typically due to systematic noise.
@@
Two challenges are scale and density variation. Nearly one million points per scan can lead to excessive memory requirements, and scanner geometry leads to uneven point densities. We partition clouds into azimuth-elevation tiles processed in batches, padded with $k_{\text{b}}$ neighbors to mitigate edge artifacts. For density variation, we use an adaptive neighborhood radius $r_{i}=\text{clamp}(\lambda d_{(k)},r_{\text{min}},r_{\text{max}})$ with $d_{(k)}$ is the $k$-th nearest neighbor distance, $\lambda$ is a scaling factor (e.g., $\lambda=1.5$), and $r_{\text{min}},r_{\text{max}}]=[0.02,0.3]$ m. Together, these strategies enable a scalable, consistent feature computation.
Detailed preprocessing steps, additional feature map examples, correlation analyses, and alternative dimensionality-reduction methods are provided in B (Table 9, Figs. 22-24).
@@ 2.2.2 Stage 2: Hybrid Annotation with Semi-Supervised and Active Learning
In Stage 2, we adopt a human-in-the-loop strategy that combines self-training and active learning to reduce manual labeling effort, as shown in Fig. 7.
First, a small subset of spherical projection image cubes—one per TLS scan—is manually annotated using Adobe Photoshop (Adobe Inc., 2024) Adobe Inc. [2024]. We use these seed labels for training an ensemble of three diverse 2D semantic segmentation models: UNet++, DeepLabV3+, and Segformer. Each model adopts a unique encoder backbone—i.e., ResNet‑34 He et al. [2016], EfficientNet‑B3 Tan and Le [2019], and MiT‑B1 Yu et al. [2023], respectively—to promote representational diversity and diminish correlated errors. UNet++ is well-suited for boundary recovery Zhou et al. [2018], DeepLabV3+ captures multi-scale context through atrous spatial pyramid pooling Chen et al. [2018], and SegFormer leverages hierarchical Transformers for long-range dependencies Xie et al. [2021].
@@
Table 3: Qualitative comparison of annotation workflows with and without uncertainty guidance.
Aspect Manual Uncertainty-assisted
Initial state Blank or uniform mask Model-generated pseudo-label
Annotation mode Labeling from scratch Correcting predictions
Edit focus Entire scene High-uncertainty regions
Error localization Visual inspection only Uncertainty map assisted
Low-uncertainty areas Manually verified Conditionally accepted
Annotator load High Reduced
Because the input feature maps extend beyond standard RGB, we re-architect each backbone into a multi-encoder fusion design (Fig. 8(a)). Each 3-channel feature group (e.g., intensity, range, Z-inv) is processed by a dedicated encoder (ResNet-34, EfficientNet-B3, or MiT-B1). These encoders are initialized with ImageNet pre-trained weights, ensuring transfer of learned low-level filters while allowing adaptation to non-RGB feature groups. Deep features are concatenated at the bottleneck and aligned by interpolation before decoding, a strategy shown to outperform naive early fusion Saidi et al. [2024], Lei et al. [2021], Zhang et al. [2020]. This modular design accommodates any $3\times N$ channel configuration with minimal retraining.
The ensemble combines predictions to generate segmentation masks and epistemic uncertainty maps: high-uncertainty regions are refined manually (active learning), while high-confidence predictions are retained as pseudo-labels (self-training). Fusing logits allows the networks to complement one another Dietterich [2000], while deep ensembles provide more reliable uncertainty estimates than a Monte-Carlo dropout approach  Lakshminarayanan et al. [2017], Kendall and Gal [2017]. The inference-refinement cycle is repeated until all scans are fully annotated. In practice, we display the pseudo-label masks, the uncertainty maps, and the input feature maps to guide annotators. As summarized in Table 3,, this joint view accelerates review by steering attention to the genuinely ambiguous regions, while leaving well-segmented areas largely untouched. Technical details of the Dice-CrossEntropy joint loss Milletari et al. [2016], Sudre et al. [2017], Mao et al. [2023a] and the ensemble-based epistemic uncertainty formulation are provided in  C (Eqs. 3-9).